The paint in a certain container is sufficient to paint an area equal to 9.375 m3• How many bricks of dimensions
22.5 cm x 10 cm x 7.5 cm can be painted out of this
contair?

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Surface area of one brick = 2(22.6 x 10 + 10 x 7.5 + 22.5 x 7.6) cm? = 2(226 + 75 + 168.76) em.2 = 937.5 cm.2• .. Number of bricks which can be painted 9.375 x 10000 = 937.5 = lOO.